OpenStudy (anonymous):

what is 95 20% off? what is 95 20% off? @Mathematics

OpenStudy (anonymous):

76

OpenStudy (anonymous):

thanks tht helped can you show me how you got the answer?

OpenStudy (anonymous):

thanks tht helped can you show me how you got the answer?

OpenStudy (anonymous):

95*(1-0.20) = 95* 0.8 = 76

OpenStudy (anonymous):

Well I'm not exactly sure how you'd set up a proportion if that's what you're looking for, but an easy way to look at it is to take 10% and double it, then subtract from 95. 10% of 95 = 9.5 9.5 x 2 = 19 95 - 19 = 76